OpenStudy (anonymous):

Marquette has $150 in his bank account at the beginnin of week 1 and will add $25 at the beginning of each subsequent week for tutoring a neighbor in math. What is the explicit formula for this sequence? Write in terms of A(n) and use no spaces in your formula.

OpenStudy (anonymous):

Week\(\color{red}{1}\), account \(A_\color{red}{1}= 150 +(\color{red}{1}-1)*25\) Week\(\color{red}{2}\), account \(A_\color{red}{2}= 150 +(\color{red}{2}-1)*25\) Week\(\color{red}{3}\), account \(A_\color{red}{3}= 150 +(\color{red}{3}-1)*25\) Week\(\color{red}{4}\), account \(A_\color{red}{4}= 150 +(\color{red}{4}-1)*25\) Can you write the formula for Week\(\color{red}{n}\) ??

OpenStudy (anonymous):

uh im not good at math

OpenStudy (anonymous):

No need to good at math!! Just observe and imitate exactly what I did. ( replace 1,2,3,4 by n) . That's it